Can you help me with this problem.
f(x)=5x+4; g(x)=7x-1
and I need to solve for (f*g)(x)
Thank you

Found 2 solutions by Alan3354, fractalier:
Answer by Alan3354(69443)

Multiply f(x) and g(x)

Answer by fractalier(6550)

(f*g)(x) = f(x) * g(x) = (5x + 4)(7x - 1) = 35x^2 + 23x - 4